# DPQuant: Efficient and Differentially-Private Model Training via Dynamic Quantization Scheduling

### v1

@@ 5.3 Dynamic Layer Selection for Quantized DP-SGD Training
$p_{i}\coloneqq\frac{\exp(-\beta R(l_{i}))}{\sum_{j=1}^{n}\exp(-\beta R(l_{j}))},\qquad\text{for }i=1,\ldots,n,$
@@
As outlined in Figure 2, to quantize $k$ out of $n$ layers at each epoch, we sample a subset without replacement according to the distribution $\{p_{i}\}$. This allows us to adaptively choose the least sensitive layers for quantization, while still randomly rotating layers with similar loss impact to minimize variance over time. This selection procedure is detailed in Appendix A.8. DPQuant provides a set of tunable parameters that govern the frequency of the analysis, as well as other privacy parameters.
@@ 5.4 Privacy Accounting
Definition 3 (Sampled Gaussian Mechanism (SGM), Mironov et al. (2019)).
@@
To protect privacy, we frame this loss computation as a Sampled Gaussian Mechanism (SGM): we draw a random subsample of $D$, clip the resulting loss value to bound sensitivity, and then add Gaussian noise of scale $\sigma$. These operations correspond to steps 4–5 of Algorithm 1.
@@
1:Input: $P$ (set of quantization policies), $B$ (batches), $R$ (iterations), $\alpha$ (EMA decay), $\mathcal{C}$ (clipping norm), $\sigma$ (noise scale), $L$ (EMA loss differences)
[…]
3:for $i=1$ to $R$ do
4:  RestoreModel()
5:  for each $(x,y)\in B$ do
6:   with no quantization do $\textsc{DPSGD-Update}(M,\,\text{loss}(M(x),y))$
7:  end for
8:  $\ell_{\text{baseline}}\leftarrow\ell_{\text{baseline}}+\frac{1}{|B|}\sum_{(x,y)\in B}\text{loss}(M(x),y)$ 1 Compute baseline loss
9:end for
10:for each $p\in P$ do
11:  $\ell[p]\leftarrow 0$
12:  for $i=1$ to $R$ do
13:   RestoreModel()
14:   for each $(x,y)\in B$ do
15:     with policy $p$ do $\textsc{DPSGD-Update}(M,\,\text{loss}(M(x),y))$
16:   end for
17:   $\ell[p]\leftarrow\ell[p]+\frac{1}{|B|}\sum_{(x,y)\in B}\text{loss}(M(x),y)$ 2 Compute loss for policy $p$
18:  end for
19:end for
@@
24:for each $p\in P$ do
25:  $L[p]\leftarrow(1-\alpha)\cdot L[p]+\alpha\cdot\hat{\mathbf{R}}[p]$ 5 Update loss difference EMA
26:end for
27:return $L$
@@
Algorithm 1 is a Sampled Gaussian Mechanism (SGM) with sample rate $q=n_{\text{sample}}/|B|$ and noise scale $\sigma=\sigma_{\text{measure}}$.
Proof: See Appendix A.4.
To account for the privacy cost we incur by performing the analysis in Algorithm 1, we rely on Opacus’s privacy accountants Yousefpour et al. (2022). This is for two reasons: First, these accountants measure the cumulative privacy loss of SGMs Makni et al. (2025), where Prop. 2 we can reuse its implementation. Second, by leveraging the advanced composition theorem Abadi et al. (2016), we obtain a much tighter upper bound on the total privacy expenditure incurred by both the DP-SGD training process and any subsequent analyses performed under the same privacy budget. We explain this in more detail in Appendix A.7.
(a) Total privacy consumption of analysis + training
(b) Fraction of privacy spent on analysis rel. training
@@ 6 Evaluation
Models and Datasets. We evaluate our approach on commonly used neural networks for differentially private training Jagielski et al. (2020); De et al. (2022): ResNet18 He et al. (2015), ResNet50 and DenseNet121 Huang et al. (2018) from in the torchvision maintainers and contributors (2016) library. These models are trained on the Extended MNIST Cohen et al. (2017), German Traffic Sign Recognition Benchmark (GTSRB) Stallkamp et al. (2011), and CIFAR-10 Krizhevsky (2009) datasets.
Hardware. The experiments were conducted on a server with a Intel(R) Xeon(R) Platinum 8470 CPU, 1TiB of memory, and 8 NVIDIA H100 80GB HBM3 GPUs.
Implementation. DPQuant is implemented on top of Opacus (Yousefpour et al., 2022), a DP training framework. Opacus provides Poisson sampling, gradient clipping, noising, and gradient accumulation.
Low Precision Format. For low precision computations, we used the LUQ-FP4Chmiel et al. (2024) format, the highest-performing 4-bit quantization format. LUQ-FP4 uses a 4-bit representation of floating point numbers. The 4-bit representation consists of 1 sign and 3 exponent bits. LUQ-FP4 first normalizes the values in a tensor, then stochastically quantizes it to one of the 16 levels.
Measurements. We use the default parameters in Table 1 unless otherwise noted. We list more detailed training parameters in Appendix A.1.
@@ 6.1 Quantization-quality trade-off
In Figure 4, we sampled $\approx 50$ random subsets of layers to execute in fp4. We plotted the empirical Pareto front using these sampled measurements, in addition to the resulting validation accuracy when using DPQuant’s scheduling technique for a given computational budget.
@@ 6.2 Sensitivity to Privacy Budget
We compared our method to the baseline for two privacy budgets $\varepsilon=4$ and $\varepsilon=8$. In Table 2 we plotted the validation accuracy for different privacy budgets. For ResNet18/50, we obtained these values by truncating the training at the respective privacy budgets (i.e. without additional hypermeter tuning), and selected baseline data point with larger $\varepsilon$ than ours wherever possible.
In most cases, DPQuant outperforms the baseline performance by at least 1 standard deviation whilst not exceeding the privacy budget. In particular, despite the privacy cost of analysis being more dominating during the $\varepsilon=4$ case, DPQuant produces near-optimal quantization schedules, demonstrating its robustness with respect to $\varepsilon$.
@@
6.3 Ablation Study
In order to better understand the contributions of the two approaches, we compared our approach (probabilistic layer sampling + loss-aware layer prioritization) with probabilistic layer sampling (PLS) alone. In Figure 5, we observe that PLS consistently performs better than the baseline where the quantized layers are selected statically. However, there is still a large gap between PLS and the best-performing layer selections, suggesting that some crucial layers are consistently being subjected to quantization which significantly degrades the quality of trained models.
When PLS is combined with loss-aware layer prioritization, the layers crucial to model training are left in full precision, even when most of the layers are quantized. The benefits of prioritization begins to surface as the proportion of quantized layers increase, as the critical layers have a larger probability of being quantized in the randomized baselines.
@@ 6.4 Theoretical Speedup
As hardware with support for FP4 MatMuls and Conv2D (e.g., NVIDIA Blackwell) are not yet widely available, we are unable to evaluate the speed benefits of quantization with DPQuant. Instead, we use estimates from prior work, along with performance statistics published by NVIDIA NVIDIA Corporation (2024) to estimate speedups. We estimate that FP4 can provide a $4\times$ speedup over the FP16 baseline by emulating FP4 computation on existing hardware. Separately, prior works Sun et al. (2020); Choi et al. (2018); Abdolrashidi et al. (2021) report a $4-7.3\times$ speedup when using FP4 on supported hardware. To remain conservative, we use the lower bound ($4\times$) in our estimates. We assume matrix multiplications, convolutions, and element-wise operations can be accelerated $4\times$, and characterize the total runtime as a linear compute cost model:
$T_{\text{ours}}=T_{\text{analysis}}+(1-p+p/4)(T_{\text{train baseline}}-T_{\text{overhead}})+T_{\text{analysis}}+T_{\text{overhead}}$
@@ 7 Conclusion
In this paper, we introduce DPQuant, a mechanism for efficient quantized DP-SGD training. We make the observation that existing quantized training techniques can significantly degrade the accuracy of models trained with DP-SGD and provided justification which demonstrated the amplified quantization error. To address this challenge, DPQuant employs techniques to dynamically select layers to quantize such that impact of quantization on model accuracy is minimized. DPQuant itself is a differentially private mechanism that incurs only small privacy cost. We empirically demonstrate that DPQuant achieves near-optimal compute-to-accuracy tradeoffs during quantized training, generalizes to different models, datasets and privacy budgets, and can provide up to $2.21\times$ speedup while minimally impacting accuracy. DPQuant enables efficient and practical differentially-private training for both centralized and distributed training deployments.
@@ Appendix A Appendix / supplemental material
A.1 Training Hyperparameters
While the learning rate might seem too high for regular SGD training, previous results Morsbach et al. [2024], Ponomareva et al. [2023] have shown that large learning rates are more beneficial for DP-SGD training.
@@ A.1 Training Hyperparameters
A.2 Accuracy Degradation for DP-SGD under Naive Quantization
Prior works Sun et al. [2020], Chmiel et al. [2024], Micikevicius et al. [2022] have demonstrated minimal degradation during quantized fp4/8 training compared to the full precision counterpart. We tabulate their results below:
@@ A.2 Accuracy Degradation for DP-SGD under Naive Quantization
Model Baseline Ultra-Low Sun et al. [2020] LUQ Chmiel et al. [2024]
@@
A.3 Proof of Proposition 1

### v2

@@ 5.3 Dynamic Layer Selection for Quantized DP Training
$\pi_{i}\coloneqq\frac{\exp(-\beta R(l_{i}))}{\sum_{j=1}^{n}\exp(-\beta R(l_{j}))},\qquad\text{for }i=1,\ldots,n,$
@@
As outlined in Figure 2, to quantize $k$ out of $n$ layers at each epoch, we sample a subset without replacement according to the distribution $\{\pi_{i}\}$. This allows us to adaptively choose the least sensitive layers for quantization, while still randomly rotating layers with similar loss impact to minimize variance over time. This selection procedure is detailed in Appendix A.15. DPQuant provides a set of tunable parameters that govern the frequency of the analysis, as well as other privacy parameters.
@@ 5.4 Privacy Accounting
Definition 3 (Sampled Gaussian Mechanism (SGM), (Mironov et al., 2019)).
@@
To protect privacy, we frame this loss computation as a Sampled Gaussian Mechanism (SGM): we draw a random subsample of $D$, clip the resulting loss value to bound sensitivity, and then add Gaussian noise of scale $\sigma$. These operations correspond to step 3 of Algorithm 1.
@@
1:Input: $P$ (policies), $B$ (batches), $R$ (iterations), $\alpha$ (decay), $\mathcal{C}$ (norm), $\sigma$ (noise)
2:Let $p_{0}$ be the baseline policy (no quantization)
3:Initialize a map for average losses, $\bar{\ell}$
4:for each $p\in P\cup\{p_{0}\}$ do $\triangleright$ 1 Compute avg. loss for baseline and all policies
5:  $\text{total\_loss}\leftarrow 0$
6:  for $i=1$ to $R$ do
7:   RestoreModel()
8:   for each $(x,y)\in B$ do
9:     With policy $p$, run $\textsc{DPSGD-Update}(M,\text{loss}(M(x),y))$
10:   end for
11:   $\text{total\_loss}\leftarrow\text{total\_loss}+\frac{1}{|B|}\sum_{(x,y)\in B}\text{loss}(M(x),y)$
12:  end for
13:  $\bar{\ell}[p]\leftarrow\text{total\_loss}/R$
14:end for
@@
19:for each $p\in P$ do
20:  $L[p]\leftarrow(1-\alpha)\cdot L[p]+\alpha\cdot\hat{\mathbf{R}}[p]$ $\triangleright$ 4 Update Policy EMA
21:end for
22:return $L$
@@
Algorithm 1 is a Sampled Gaussian Mechanism (SGM) with sample rate $q=|B|/|D|$ and noise scale $\sigma=\sigma_{\text{measure}}$. Proof: See Appendix A.11.
To account for the privacy cost we incur by performing the analysis in Algorithm 1, we rely on Opacus’s privacy accountants (Yousefpour et al., 2022). This is for two reasons: First, these accountants measure the cumulative privacy loss of SGMs (Makni et al., 2025), where by Prop. 2 we can reuse its implementation. Second, by leveraging the advanced composition theorem (Abadi et al., 2016), we obtain a much tighter upper bound on the total privacy expenditure incurred by both the DP-SGD training process and any subsequent analyses performed under the same privacy budget. We explain this in more detail in Appendix A.14.
(a) privacy consumption of analysis + training
(b) Fraction of privacy spent on analysis
@@ 6 Evaluation
Models and Datasets. We evaluate our approach on commonly used neural networks for differentially private training (Jagielski et al., 2020, De et al., 2022): ResNet18 (He et al., 2015), ResNet50 and DenseNet121 (Huang et al., 2018) from the torchvision (maintainers and contributors, 2016) library. We also test BERT (Devlin et al., 2019), trained with DP-AdamW. We additionally evaluate with DP-Adam in Appendix A.5. These models are trained on the Extended MNIST (Cohen et al., 2017), German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2011), CIFAR-10 (Krizhevsky, 2009) and SNLI (Bowman et al., 2015) datasets.
Implementation. DPQuant is implemented on top of Opacus (Yousefpour et al., 2022), a DP training framework which provides Poisson sampling, gradient clipping, and noising. The DPQuant parameters can be found in Appendix A.2.
Low Precision Format. For low precision computations, we used the LUQ-FP4 (Chmiel et al., 2024) format, the highest-performing 4-bit quantization format. LUQ-FP4 uses a 4-bit representation of floating point numbers, consisting of 1 sign and 3 exponent bits. In Appendix A.9, we evaluate DPQuant on other low-precision formats including FP8 and 4-bit uniform quantization.
@@ 6.1 Quantization-quality trade-off
In Figure 4, we sample $\approx 50$ random subsets of layers to execute in fp4. We plot the empirical Pareto front using these sampled measurements, in addition to the resulting validation accuracy when using DPQuant’s scheduling technique for a given computational budget.
@@ 6.2 Sensitivity to Privacy Budget
BERT SNLI 0.5 $62.54\text{\,}$$\pm$$4.54\text{\,}$ $7.48\text{\,}$ $67.80\text{\,}$ $7.48\text{\,}$
0.75 $52.04\text{\,}$$\pm$$3.95\text{\,}$ $7.48\text{\,}$ $63.61\text{\,}$ $7.48\text{\,}$
@@
We compared our method to the baseline for two privacy budgets $\varepsilon=4$ and $\varepsilon=8$. In Table 1 we report the validation accuracy for different privacy budgets. For ResNet18/50, we obtain these values by truncating the training at the respective privacy budgets (i.e. without additional hyperparameter tuning), and select baseline data points with larger $\varepsilon$ than ours wherever possible.
In most cases, DPQuant outperforms the baseline performance by at least 1 standard deviation whilst not exceeding the privacy budget. In particular, despite the privacy cost of analysis being more dominant during the $\varepsilon=4$ case, DPQuant produces near-optimal quantization schedules, demonstrating its robustness with respect to $\varepsilon$.
We have also evaluated DPQuant on extremely small privacy budgets (e.g. $\varepsilon=1$). In Appendix A.3 we show that DPQuant still demonstrates the same benefits under this setting.
6.3 Ablation Study
To better understand the contributions of the two approaches, we compared our approach (probabilistic layer sampling + loss-aware layer prioritization) with probabilistic layer sampling (PLS) alone. In Figure 5, we observe that PLS consistently performs better than the baseline where the quantized layers are selected statically. However, there is still a large gap between PLS and the best-performing layer selections, suggesting that some crucial layers are consistently being subjected to quantization which significantly degrades the quality of trained models.
When PLS is combined with loss-aware layer prioritization, the layers crucial to model training are left in full precision, even when most of the layers are quantized. The benefits of prioritization begins to surface as the proportion of quantized layers increase, as the critical layers have a larger probability of being quantized in the randomized baselines. Furthermore, we observe that the best training outcomes are achieved by combining both approaches. We include more details in Appendix A.7.
@@ 6.4 Theoretical Speedup
As hardware with support for FP4 MatMuls and Conv2D (e.g., NVIDIA Blackwell) are not yet widely available, we are unable to evaluate the speed benefits of quantization with DPQuant. Instead, we use estimates from prior work, along with performance statistics published by NVIDIA (NVIDIA Corporation, 2024) to estimate speedups. Based on profiling on existing hardware, we estimate that FP4 can provide a $4\times$ speedup over the FP16 baseline. Separately, prior works (Sun et al., 2020, Choi et al., 2018, Abdolrashidi et al., 2021) report a $4-7.3\times$ speedup when using FP4 on supported hardware. To remain conservative, we use the lower bound ($4\times$) in our estimates. We assume matrix multiplications, convolutions, and element-wise operations can be accelerated $4\times$, and characterize the total runtime as a linear compute cost model:
$T_{\text{ours}}=T_{\text{analysis}}+(1-p+p/4)(T_{\text{train baseline}}-T_{\text{overhead}})+T_{\text{overhead}}$
@@ 7 Conclusion
In this paper, we introduce DPQuant, a mechanism for efficient quantized DP training. We observe that existing quantized training techniques can significantly degrade the accuracy of models trained with DP-SGD and DP-Adam, and provide a detailed analysis demonstrating the amplified quantization error. To address this challenge, DPQuant employs techniques to dynamically select layers to quantize such that the impact of quantization on model accuracy is minimized. DPQuant itself is a differentially private mechanism that incurs only small privacy cost. We empirically demonstrate that DPQuant achieves near-optimal compute-to-accuracy tradeoffs during quantized training, generalizes to different models, datasets and privacy budgets, and can provide up to $2.21\times$ speedup while minimally impacting accuracy. DPQuant enables efficient and practical differentially-private training for both centralized and distributed training deployments.
@@ Appendix A Appendix / supplemental material
A.1 Effect of Batch Size
Our analysis of quantization error in DP-SGD is independent of batch size. While larger batches reduce stochastic gradient variance, our argument hinges on the magnitude of the final noised gradient, which remains large regardless of batch size.
  1. 1.
    In DP-SGD, the added noise’s scale is proportional to the per-example clipping constant, $C$, not the batch size.
  2. 2.
    This noise dominates the averaged gradient signal, causing the raw gradients in subsequent steps to have significantly larger norms than in non-DP training.
  3. 3.
[…]
To demonstrate this empirically, we ran the same training job with batch sizes ranging from 1024 to 8192 and measured the numerical range of the weight gradients, similar to that in figure 1c. Across the batch sizes, there is negligible difference in the gradient ranges, which confirms our hypothesis.
Table 2: Weight gradient norm range across various batch sizes, showing the negligible impact of batch size on the final gradient magnitudes.
Batch Size Norm Range Mean Norm Range Std
1024 0.159 0.137
2048 0.161 0.127
4096 0.158 0.116
8192 0.156 0.119
A.2 Evaluation Setup and Parameters
Parameter Default Description
$n$ 60 number of epochs to train
$k$ layers to execute in low precision.
$n_{\text{sample}}$ 1 test samples for loss measurement.
$n_{\text{interval}}$ 2 epochs to train before the next measurement.
$R$ 2 repetitions during measurement.
$\sigma_{\text{measure}}$ 0.5 Noise scale used during loss‐difference privatization.
$\mathcal{C}_{\text{measure}}$ 0.01 Clipping norm used during loss‐difference privatization.
Table 3: Configurable Hyperparameters of DPQuant
Remark: Selecting DPQuant parameters in practice.
In our experiments, we have found repetitions = 2 and sampling frequency = 1 to be the most optimal. Adopting these recommended defaults, the user needs to pick:
  1. 1.
    one of $k$ (number of layers to quantize) and the analysis frequency
  2. 2.
    clipping norm used in loss sensitivity analysis
The process of determining clipping norm for analysis is similar to that of finding the clipping threshold C for normal DP-SGD training. We want to pick a value such that the differences between policies are expressed.
A.3 Evaluation under Extreme Privacy Budgets
As shown in Figure 3, the privacy consumption of DPQuant’s analysis accounts for a higher fraction of the total privacy near the beginning of training. We wish to evaluate DPQuant under more strict privacy budgets.
In these cases, both the parameters of DP-SGD and DPQuant need to be updated, namely the noise scale $\sigma$ and measurement noise scale $\sigma_{\text{measure}}$ both need to be increased. Table 4 below shows that DPQuant achieves optimal accuracy even when $\varepsilon=1$.
Table 4: Training accuracy of ResNet18 with GTSRB under the strict privacy budget ($\varepsilon=1$)
Baseline DPQuant
Count Accuracy (%) $\varepsilon$ Accuracy (%) $\varepsilon$
50% 44.14 $\pm$ 4.61 1.05 48.26 0.99
75% 40.13 $\pm$ 3.65 1.05 43.14 0.99
90% 35.20 $\pm$ 1.12 1.05 38.66 0.99
A.4 Training Hyperparameters
A.4.1 Image Models
While the learning rate might seem too high for regular SGD training, previous results Morsbach et al. (2024), Ponomareva et al. (2023) have shown that large learning rates are more beneficial for DP-SGD training.
@@ A.4.1 Image Models
A.4.2 Language Models
We conducted a new NLP experiment using BERT for sequence classification on the Stanford Natural Language Inference (SNLI) corpus. In this task, the model classifies a pair of statements (e.g., “Children smiling and waving at camera” and “There are children present”) as “entailment,” “contradiction,” or “neutral.”
Due to the high number of parameters in BERT, we have followed the tutorial from Opacus and frozen 12 out of 13 BERT layers, and trained the last BERT layer and subsequent classification layers.
We have compared our method (DPQuant) with a random static baseline (similar to section 6.1). We use the same training parameters, trained for a single epoch, and used $\varepsilon=8$ as the total privacy budget.
In these experiments, DPQuant outperforms the baseline in accuracy. DPQuant consistently avoids quantizing the last few layers (including the trainable ones) without prior information about the importance and trainability of the layers.
A.5 Evaluation with DP-Adam
We evaluate DPQuant using DP-Adam as the optimizer. All hyperparameters match the DP-SGD configuration in Table 5, with the optimizer changed to Adam (learning rate $0.01$, default $\beta_{1}=0.9$, $\beta_{2}=0.999$).
Table 6 shows that DPQuant consistently outperforms the static baseline under DP-Adam while consuming a strictly lower privacy budget, with the largest gains at 75% and 90% quantization.
Table 6: DP-Adam: DPQuant vs. static random baseline (ResNet18).
Baseline DPQuant
Dataset Percent Quantized Accuracy (%) $\varepsilon$ Accuracy (%) $\varepsilon$
CIFAR-10 0.50 $54.18\pm 1.48$ 5.05 55.03 4.97
0.75 $49.86\pm 1.57$ 5.05 52.40 4.97
0.90 $47.36\pm 1.36$ 5.05 51.18 4.97
GTSRB 0.50 $72.74\pm 6.04$ 6.05 75.87 5.91
0.75 $58.14\pm 7.64$ 6.05 72.32 5.91
0.90 $50.88\pm 5.11$ 6.05 66.45 5.91
Additionally, the BERT experiment on SNLI (Table 1) uses DP-AdamW as the underlying optimizer.
A.6 Accuracy Degradation for DP-SGD under Naive Quantization
Prior works Sun et al. (2020), Chmiel et al. (2024), Micikevicius et al. (2022) have demonstrated minimal degradation during quantized fp4/8 training compared to the full precision counterpart. We tabulate their results below:
@@ A.6 Accuracy Degradation for DP-SGD under Naive Quantization
Model Baseline Ultra-Low Sun et al. (2020) LUQ Chmiel et al. (2024)
@@
A.7 Sensitivity of temperature $\beta$
In our method, the temperature parameter $\beta$ provides a crucial mechanism to balance two complementary strategies:
  • Deterministic Selection: This approach prioritizes elements based on their loss sensitivity, selecting those that are most impactful.
  • Randomized Sampling: This approach introduces stochasticity, ensuring diversity and exploration in the selection process.
A low $\beta$ value favors randomized sampling, while a high $\beta$ value makes the selection process more deterministic and reliant on loss sensitivity. Below, we tabulate the training accuracy for different value of $\beta$, and observe that better training outcomes can be obtained by favoring loss-based layer selection while retaining some stochasticity. Namely, it performs strictly better than selection purely based on random layer sampling.
Table 9: Model performance across various counts and temperature ($\beta$) values
Temperature ($\beta$)
Count 0.1 0.22 0.47 1.03 2.24 4.86 10.57 22.99 50.0
10 $66.49$ $67.58$ $67.53$ $67.01$ $70.25$ $70.37$ 71.59 $70.96$ 71.67
15 $58.47$ $58.47$ $60.03$ $59.07$ $60.86$ 65.00 $60.54$ 65.04 $63.75$
18 $51.60$ $54.08$ $55.73$ $53.73$ $53.86$ $53.49$ 60.90 $55.45$ $56.05$
A.8 EMA Ablation
The exponential moving average (EMA) in step  4 of Algorithm 1 smooths the per-layer importance scores across measurement iterations so that a single noisy batch does not drastically change the layer ranking. Table 10 shows the effect of removing EMA on ResNet18 trained on CIFAR-10 with DP-SGD. Across all quantization levels, EMA consistently improves accuracy, confirming that smoothing the noisy sensitivity estimates is beneficial.
Table 10: Effect of EMA on DPQuant (ResNet18, CIFAR-10, DP-SGD).
Percent Quantized With EMA Without EMA
0.50 56.06 54.01
0.75 51.59 50.07
0.90 49.97 47.22
A.9 Evaluation on other Quantizers
To assess the versatility of our method, DPQuant, we evaluated its performance with different numerical precisions and quantization schemes. We conducted two primary experiments: one using 8-bit floating-point (fp8e5m2) for training to test a different bitwidth, and another using a uniform 4-bit quantizer to test a different quantization strategy.
A.9.1 FP8 Quantization
Our experiments with FP8 training show that quantized DP-SGD does not suffer a significant performance degradation. This minimal performance gap suggests that in higher-precision settings like FP8, the benefits of more complex techniques like layer subset selection may be less critical. We show the results in table 11.
Table 11: Performance comparison with FP8 training.
Count Base Acc(%) Base $\epsilon$ Our Acc(%) Our $\epsilon$
50% $67.56\pm 0.47$ 4.05 67.12 3.93
75% $67.76\pm 0.67$ 4.05 67.65 3.93
90% $67.38\pm 0.59$ 4.05 68.01 3.93
A.9.2 Uniform 4-bit Quantization
Next, we evaluated a more aggressive quantization scheme using a uniform FP4 quantizer. In this setup, the value range is discretized into $2^{4}=16$ levels via stochastic rounding. The results reveal a more substantial drop in accuracy for our method compared to the baseline. This outcome is consistent with our observations of the LUQ-FP4 quantizer discussed in Section 6.2, highlighting the inherent challenges of applying DP-SGD with very low-bitwidth uniform quantization. We show the results in table 12.
Table 12: Performance comparison with uniform FP4 quantization.
Count Base Acc(%) Base $\epsilon$ Our Acc(%) Our $\epsilon$
50% $63.56\pm 0.89$ 4.53 62.15 4.44
75% $57.85\pm 0.90$ 4.53 59.09 4.44
90% $55.82\pm 0.80$ 4.53 56.27 4.44
A.10 Proof of Proposition 1